# 3D-Prover: Diversity Driven Theorem Proving With Determinantal Point Processes

### v1

@@ 3.2 Filtering Model
1 for $i$ in $\{1,\dots,N\}$ do
       $\bm{\phi}_{i}\leftarrow E(g,t_{i})$ ;
        // Compute tactic embedding
       $(\tau_{i},s_{i})\leftarrow P(\bm{\phi}_{i})$ ;
        // Compute time and error scores
       $\tau_{i}\leftarrow 1-\frac{\tau_{i}}{||\bm{\tau}||},\bm{\tau}=(\tau_{1},..,\tau_{N})$ ;
        // Normalise time scores
       $m_{i}\leftarrow\frac{\text{exp}(\pi_{0}(t_{i}|g)\mathbin{/}\theta)}{\sum_{j=1}^{N}\text{exp}(\pi_{0}(t_{j}|g)\mathbin{/}\theta)}$ ;
        // Normalise model logits
       $q_{i}\leftarrow m_{i}+\lambda_{s}s_{i}+\lambda_{\tau}\tau_{i}$ ;
        // Compute quality score
2
// Filter tactics with $k$-DPP
$L\leftarrow B^{T}B$, where $B=[q_{1}\bm{\phi}_{1},\dots,q_{N}\bm{\phi}_{N}]$ ;
@@
3 Compute eigenvalues $\lambda_{i}$ and eigenvectors ${\bm{v}}_{i}$ of $L$
4 Sample $J\subset\{1,\dots,N\}$ using Algorithm 2 of Kulesza & Taskar (2011),
5   with parameters $\{({\bm{v}}_{i},\lambda_{i})\}$, $k=K$
6 return $T^{\prime}=\{t_{j}\}_{j\in J}$
@@ 4 Sample $J\subset\{1,\dots,N\}$ using Algorithm 2 of Kulesza & Taskar (2011),
Algorithm 1 defines our filtering model, 3D-Prover, which maps a list of tactics $T$ from the underlying tactic policy $\pi_{0}$ to a subset $T^{\prime}$ of size $K$. We use the Encoder $E$ and Predictor $P$ defined in Section 1.2 to generate unit norm tactic embeddings $\bm{\phi}_{i}$ and predict the time and error likelihood. The embeddings $\bm{\phi}_{i}$ encode the predicted environment response through their direction only, as they are unit norm (Figure 3). The quality scores $q_{i}$ then scale these tactics based on the underlying model logits $m_{i}$, as well as the predicted error likelihood $s_{i}$ and execution time $\tau_{i}$. We have hyperparameters for the normalisation temperature $\theta$, as well as the error and time weights $\lambda_{s},\lambda_{\tau}$. The parameter $\theta$ controls the scaling temperature of the model logits, with a higher temperature flattening out the distribution. It therefore adjusts the diversity bias of the filtering model by reducing the impact of the quality scores when sampling. We then compute the kernel $L$ from $q_{i}$ and $\bm{\phi}_{i}$, and sample a subset of tactics $T^{\prime}$ using the $k$-DPP algorithm (Kulesza & Taskar, 2011).
@@ 3.3 Experiments
To test the performance of 3D-Prover, we use ReProver (Yang et al., 2023) as the underlying tactic policy $\pi_{0}$, with the Encoder $E$ and Predictor $P$ components as defined in Section 1.2. We chose ReProver as it is a small ($\sim$ 300M parameters), popular and performant open source model, allowing us to run our experiments in a reasonable timeframe. We run our experiments in Lean 3 (De Moura et al., 2015) using the BAIT (Lamont et al., 2024) platform with a modified LeanDojo (Yang et al., 2023) environment, where we set an environment timeout of 600 seconds per proof attempt. We train a combined transition model on the miniF2F-valid benchmark, and use the Encoder and Predictor components to generate tactic embeddings and quality scores as per Algorithm 1. We first examine the performance of 3D-Prover without any hyperparameter tuning, setting $\lambda_{s}=\lambda_{\tau}=0$, $\theta=1$. We then perform ablation studies using miniF2F-valid to examine the influence of the hyperparameters on the tactic success rate, execution time and diversity of the environment response. For miniF2F-test, we allow the model four attempts per proof to increase confidence in the results, while for miniF2F-valid we allow one attempt per configuration to facilitate a wider set of ablations.
We set the search policy for all experiments to be Best First Search (BestFS), where nodes are expanded in order of their cumulative log probability. For each node selected for expansion, we generate $N=64$ candidate tactics from the underlying ReProver model using beam search with default settings, as done in the original ReProver implementation. This forms the ground set for the node, to be sub-sampled by the filtering algorithm. We use beam search decoding because it is deterministic and so ensures that the ground set for a given node remains fixed across runs, allowing us to isolate and compare the effect of the filtering algorithm. The filtering algorithm returns $K$ tactics, which are then executed in the environment and used to update the proof tree, as outlined in 1.2. We test three different levels of filtering, with $K\in\{8,16,32\}$. Lower values of $K$ correspond to more filtering, for which the choice of filtering algorithm will have a greater impact. We compare the filtering approach of 3D-Prover, as outlined in Algorithm 1, with two baselines. The Top-$\mathbf{K}$ baseline takes the top $K$ tactics from the ground set as judged by their log probabilities, corresponding to the top $K$ beams. We take $K$ tactics at random from the ground set to form the Random baseline, as an exploration-focused comparison.
@@ 3.3.1 Proof Performance
miniF2F-test miniF2F-valid
$K$ Top-$K$ Random 3D-Prover Gain Top-$K$ Random 3D-Prover Gain
8 22.4 19.0 $\pm$ 0.98 24.4 $\pm$ 0.22 +8.9% 21.7 19.3 25.0 +15.2%
16 26.5 25.4 $\pm$ 0.39 27.3 $\pm$ 0.21 +3.0% 26.6 24.2 29.1 +9.4%
32 27.8 27.4 $\pm$ 0.26 28.2 $\pm$ 0.25 +1.4% 27.9 27.5 28.7 +2.9%
@@
Table 2: Percentage of proofs found after one attempt (Pass@1) on miniF2F, with $K$ tactics selected per node, using tactics generated from ReProver. 3D-Prover uses a transition model trained from miniF2F-valid transitions. For miniF2F-test, we report the mean $\pm$ standard error over four runs, with Top-$K$ being deterministic. The Gain column reports the relative improvement over the Top-$K$ baseline. We observe a clear improvement using 3D-Prover, which increases as more filtering is applied (lower $K$). Our results on miniF2F-test show that 3D-Prover can improve search even for proofs out of distribution of the transition model.
@@
Table  2 shows the Pass@1 results of our experiments on miniF2F, which is the number of proofs successfully found after a single attempt. We observe that 3D-Prover significantly outperforms both baseline approaches. We also note that Top-$K$ selection performs better than the Random approach, which is unsurprising. The influence of the filtering algorithm becomes more apparent as $K$ is decreased, as there are more tactics filtered out. Our results are consistent with this, with the magnitude of improvement given by 3D-Prover increasing for lower values of $K$. 3D-Prover is able to outperform both baselines by providing a tradeoff between the quality, as represented by Top-$K$, and the diversity of the tactics. The choice of $K$ also controls the depth of the proof search, with larger $K$ giving broader search, and smaller $K$ deeper search. As most discovered proofs are short (favouring broad search), the Pass@1 performance for lower values of $K$ is generally lower, however over multiple attempts it can be beneficial to use deeper searches (see Appendix A.1). Finding deep proofs has to date been a significant challenge ( e.g. Polu et al. (2022)), with the search tree growing exponentially with the proof depth. The improvement given by 3D-Prover, particularly for deeper search configurations, is a step towards addressing this.
@@
Tree search should be considered as an augmentation of the base model, with the degree of any improvement much smaller than what can be found by improving the generator itself. This is unsurprising, as the generator forms the base set of candidates for the search to explore. Improved search algorithms do however have the advantage of being applicable to different base models, which is important given the rapid advancement of new and better generators. For example, the state-of-the-art DeepSeek-Prover-V1.5 (Xin et al., 2024) obtains around 2–4% relative improvements in proof success over miniF2F-test with its novel tree search algorithm, compared to no search (Figure 5). In comparison, improving their base model yields a $\sim$36relative improvement (Table 1). Similarly, Han et al. (2022) obtain $\sim$6–9relative improvements (Table 7) for search, with $\sim$40,000 GPU hours required for their best results. We were able to find our improvements with significantly less resources, training our transition model on only a single attempt per proof.
@@
We emphasise that these results were obtained without any hyperparameter tuning, only using the representations as diversity features and model logits as quality scores. We present ablation studies looking closer at these hyperparameters, however a comprehensive sweep is prohibitively expensive with each full attempt taking at least 12h on our hardware. Despite this, we were able to obtain our improvements without any tuning, demonstrating the effectiveness of our approach. We also highlight that the miniF2F-test results were obtained by training with transitions from miniF2F-valid, showing that 3D-Prover remains effective for proofs out of distribution.
@@ 3.3.2 Ablation Study
$K$ Autoencoder Transition Model Gain
8 23.0 25.0 +8.7%
16 27.9 29.1 +4.3%
32 27.0 28.7 +6.3%
@@
To demonstrate the utility of our transition model representations, we compare to an ablated 3D-Prover where the transition model Encoder is replaced by an Autoencoder of the same size. The Autoencoder is trained to reconstruct the original tactic, and therefore generates representations which reflect only the syntax of the tactic. In this way, we can test our hypothesis that semantically aware tactic representations are useful for proofs, justifying the inclusion of the transition model. As we observe in Table 3, the performance of 3D-Prover with the transition model embeddings is indeed superior to that of the Autoencoder across all values of $K$. This shows that selecting for diversity with respect to the predicted semantics, rather than the syntax, leads to a direct improvement in proof performance.
3D-Prover
$K$ Top-$K$ Random $\lambda_{s}=0.1$ $\lambda_{s}=0.5$
@@
Table 4: Tactic success percentage per node for miniF2F-valid (Mean $\pm$ Standard Error), where $\lambda_{s}$ controls the error weight of quality score in 3D-Prover. These results demonstrate that 3D-Prover leads to fewer errors on average, which can be controlled by increasing $\lambda_{s}$.
Success Rate
We observe from Table 4 that the success rate of tactics chosen by 3D-Prover is significantly improved compared to both baselines. We also note that as $K$ decreases, this improvement increases in magnitude, reflecting the heightened influence of the filtering model. We see that this improvement increases with the error weight $\lambda_{s}$, which scales the quality scores of tactics by their predicted probability of success. This suggests the error weight term is directly influencing the tactic success rate, showing that it is working as intended.

### v2

@@ 2.2 Experiments
for $i$ in $\{1,\dots,N\}$ do
 $\bm{\phi}_{i}\leftarrow E(g,t_{i})$
 $(s_{i},\tau_{i})\leftarrow P(\bm{\phi}_{i})$
[…]
 $m_{i}\leftarrow\frac{\text{exp}(\pi_{0}(t_{i}|g)\mathbin{/}\theta)}{\sum_{j=1}^{N}\text{exp}(\pi_{0}(t_{j}|g)\mathbin{/}\theta)}$
 $q_{i}\leftarrow m_{i}+\lambda_{s}s_{i}+\lambda_{\tau}\tau_{i}$
@@
$L\leftarrow B^{T}B$, where $B=[q_{1}\bm{\phi}_{1},\dots,q_{N}\bm{\phi}_{N}]$
Compute eigenvalues $\lambda_{i}$ and eigenvectors ${\bm{v}}_{i}$ of $L$
Sample $J\subset\{1,\dots,N\}$ using Algorithm 2 of Kulesza & Taskar (2011), with $\{({\bm{v}}_{i},\lambda_{i})\}$, $k=K$
return $T^{\prime}=\{t_{j}\}_{j\in J}$
@@
Algorithm 1 defines 3D-Prover, which maps tactics $T$ from the underlying tactic policy $\pi_{0}$ to a subset $T^{\prime}$ of size $K$. We use the Encoder $E$ and Predictor $P$ from Section 2 to generate tactic embeddings $\bm{\phi}_{i}$ and predict the time and error likelihood. As they are unit norm, the embeddings $\bm{\phi}_{i}$ encode the predicted environment response through their direction. The quality score $q_{i}$ then scales $\bm{\phi}_{i}$ based on the tactic model logits $m_{i}$, as well as the predicted error likelihood $s_{i}$ and execution time $\tau_{i}$. We have hyperparameters for normalisation temperature $\theta$, as well as error and time weights ($\lambda_{s},\lambda_{\tau}$). $\theta$ controls the scaling temperature of the model logits, with a higher temperature flattening the distribution. It therefore adjusts the diversity bias of 3D-Prover by reducing the impact of the quality scores when sampling. We then compute the kernel $L$ from $q_{i}$ and $\bm{\phi}_{i}$, and sample a subset of tactics $T^{\prime}$ using the $k$-DPP algorithm (Kulesza & Taskar, 2011). Figure 4(a) visualises this process, where tactics subsets are sampled in proportion to their shaded area.
@@ 3.1 Experiments
We test the performance of 3D-Prover with two setups. We first use ReProver (Yang et al., 2023) as the underlying tactic policy $\pi_{0}$, as it is a small ($\sim$ 300M parameters) and popular open source proving model, allowing us to run extensive experiments and ablations in a reasonable time frame. To evaluate our approach over a large (7B) state-of-the-art model, we also present a smaller scale experiment using InternLM2.5-Step-Prover (Wu et al., 2024). We run our experiments in Lean (De Moura et al., 2015) using the BAIT (Lamont et al., 2024) platform with a modified LeanDojo (Yang et al., 2023) environment, where we set an environment timeout of 600 seconds per proof attempt. We train a Combined model for Reprover and a Combined (Small Goal) model for InternLM from Section 2.2, using transitions from their respective base models. The Encoder and Predictor components then generate tactic embeddings and quality scores as per Algorithm 1. We first examine the performance of 3D-Prover without hyperparameter tuning, setting $\lambda_{s}=\lambda_{\tau}=0$, $\theta=1$. We then perform ablation studies ( 3.1.2) with ReProver over miniF2F-valid to examine the influence of the hyperparameters on the tactic success rate, execution time and diversity of the environment response. For miniF2F-test, we allow the model multiple attempts per proof to increase confidence in the results, while for miniF2F-valid we allow one attempt per configuration to allow a wider set of ablations. We also present an additional experiment over the larger LeanDojo benchmark in Appendix D.
The default search policy is set to be Best First Search (BFS), with nodes expanded in order of their cumulative log probability. For each node selected, we generate $N=64$ candidate tactics from the underlying model  [^2] As InternLM is sampling, this is up to 64. See Appendix H for the distribution of tactic counts per node. Following the original implementations, we use beam search for ReProver and sampling with $T=0.7$ for InternLM. These form the ground set for the node, to be sub-sampled by the filtering algorithm. As beam search decoding is deterministic, the ground set for a given node is fixed across runs, allowing us to better isolate and compare approaches. We maintain sampling for InternLM to represent the original deplyoment scenario and test our approach under realistic usage. The filtering algorithm returns $K$ tactics, which are executed in the environment before updating the proof tree. For ReProver, we test three levels of filtering, with $K\in\{8,16,32\}$. Lower $K$ corresponds to more filtering, for which the filtering algorithm will have a greater impact. We compare 3D-Prover, as outlined in Algorithm 1, with three baselines. The No Filtering baseline represents the original approach with no filtering. The Top-$\mathbf{K}$ baseline takes the top $K$ tactics from the ground set as judged by their log probabilities, corresponding to the top $K$ beams. We take $K$ tactics at random from the ground set to form the Random baseline, as an exploration-focused comparison. For InternLM, we test $K=8$ with the Top-$K$ and No Filtering baselines, and perform an additional experiment with Critic Guided search from (Wu et al., 2024) in place of BFS.
@@ 3.1.1 Proof Performance
Table 2: Pass@1 results on miniF2F, with $K$ tactics selected per node from ReProver. 3D-Prover uses a transition model trained from miniF2F-valid transitions. For miniF2F-test, we report the mean along with minumum and maximum over four runs, noting that Top-$K$ is deterministic given ReProver uses Beam Search. The Gain column reports the relative improvement over the Top-$K$ baseline. Results for No Filtering were 27.8% for miniF2F-test and 27.9% for miniF2F-valid. We observe a clear improvement using 3D-Prover, which increases as more filtering is applied (lower $K$). Our results on miniF2F-test show that 3D-Prover can improve search even for proofs out of distribution of the transition model.
$K$ Top-$K$ Random 3D-Prover Gain
miniF2F-test (mean, minimum and maximum over four runs)
8 22.4 19.0 (18.4, 19.6) 24.4 (23.7, 24.9) +8.9%
16 26.5 25.4 (24.5, 25.7) 27.3 (26.9, 27.8) +3.0%
32 27.8 27.4 (26.9, 28.2) 28.2 (27.3, 28.6) +1.4%
miniF2F-valid (single run)
8 21.7 19.3 25.0 +15.2%
16 26.6 24.2 29.1 +9.4%
32 27.9 27.5 28.7 +2.9%
Table 3: Results for miniF2F-test using tactics selected from InternLM2.5-Step-Prover (mean, maximum and minimum for Pass@1). We compare results using standard Best First Search model (BFS) and with the InternLM2.5 Critic Guided model for goal selection. 3D-Prover uses a transition model trained from transitions on miniF2F-test with the No Filtering model. The Pass@1 result for BFS for No Filtering was 44.8. We observe 3D-Prover outperforming both the No Filtering and Top-$K$ baselines, with a notable improvement over more attempts for the Critic Guided search model.
Top-$K$ ($K=8)$ 3D-Prover ($K=8$) Gain
InternLM2.5-StepProver (BFS)
Pass@1 44.3 (44.1, 44.5) 45.7 (45.3, 46.1) +3.2%
Pass@2 44.9 47.3 +5.3%
No Filtering ($K=64$) 3D-Prover ($K=8$) Gain
InternLM2.5-StepProver (Critic Guided)
Pass@1 43.7 (42.4, 44.5) 45.7 (44.5, 47.0) +4.6%
Pass@6 49.0 53.1 +8.4%
@@
Table 2 and 3 shows the results of our ReProver and InternLM experiments on miniF2F. We observe 3D-Prover outperforming every baseline over both models. The influence of filtering becomes more apparent as $K$ is decreased as there are more tactics filtered out. Reflecting this, the magnitude of improvement given by 3D-Prover increases for lower $K$. 3D-Prover is able to outperform both baselines by providing a tradeoff between the quality, as represented by Top-$K$, and the diversity of the tactics. The choice of $K$ also controls the depth of the proof search, with larger $K$ giving broader search, and smaller $K$ deeper search. As most discovered proofs are short (favouring broad search), we see the Pass@1 performance for lower values of $K$ is generally lower, however over multiple attempts it can be beneficial to use deeper searches (see Appendix B). Our InternLM results (Table 3) demonstrate this, with the improvements from filtering growing over more attempts. This indicates a better variety of paths being explored, as each different attempt is more likely to take a more diverse approach. Finding deep proofs is a significant challenge Polu et al. (2022), with the search tree growing exponentially with proof depth. The large improvements from 3D-Prover for deeper search is a step towards addressing this.
@@
Tree search should be considered an augmentation of the base model, with improvements generally much smaller than those found by improving the generator itself. This is unsurprising, as the generator forms the base set of candidates for the search to explore. Improved search algorithms do, however, have the advantage of being applicable to different base models, which is important given the rapid advance of new and better generators. For example, DeepSeek-Prover-V1.5 obtains around 2–4% relative improvements in proof success over miniF2F-test with its novel tree search algorithm, compared to no search. In comparison, improving their base model yields a $\sim$36% relative improvement (Figure 5 and Table 1 in (Xin et al., 2024)). Similarly, Table 1 from Polu et al. (2022) shows their search approach yielding 0.04-5.7% relative improvements for miniF2F-valid, with $\sim$40,000 GPU hours required for their best results. We were able to find our improvements with significantly less resources, training our transition model on only a single attempt per proof.
@@
We emphasise that these results were obtained without any hyperparameter tuning, only using the representations as diversity features and model logits as quality scores. We present ablation studies looking closer at these hyperparameters, however a comprehensive sweep is prohibitively expensive (Appendix  I). Despite this, we were able to obtain our improvements without tuning, demonstrating the effectiveness of our approach. For completeness, Appendix C details the Pass@1 performance over the hyperparameter configurations we tested for our ablations. We also highlight that the miniF2F-test ReProver results were obtained by training with transitions from miniF2F-valid, showing that 3D-Prover remains effective for proofs out of distribution. The results on miniF2F-valid represent the more common online scenario, with previous attempts on the same dataset being used to improve performance (see, for example, Lample et al. (2022); Polu et al. (2022)). We also note that our approach is lightweight with minimal overhead, as we detail in Appendix G.
@@ 3.1.2 Ablation Study
To demonstrate the utility of our tactic representations, we compare to an ablated 3D-Prover where the transition model Encoder is replaced by an Autoencoder of the same size. The Autoencoder is trained to reconstruct the original tactic, and therefore generates representations which reflect only the syntax of the tactic. This tests our hypothesis that semantically aware tactic representations are useful for proofs, justifying the inclusion of the transition model. From Table 4, the performance of 3D-Prover with the transition model embeddings is indeed superior to that of the Autoencoder across all values of $K$. This shows that selecting for diversity with respect to the predicted semantics, rather than the syntax, leads to a direct improvement in proof performance.
@@
$K=8$ $K=16$ $K=32$
Autoencoder 23.0 27.9 27.0
3D-Prover 25.0 29.1 28.7
We have demonstrated that 3D-Prover improves proof success rate without any hyperparameter tuning, with a fixed $\lambda_{s}=\lambda_{\tau}=0$, $\theta=1$. We now examine whether we can use 3D-Prover to direct search to optimise secondary objectives, namely the execution time, tactic success rate and the diversity of environment response.
Success Rate
We observe from Table 5 3D-Prover significantly improves the success rate of chosen tactics. As $K$ decreases, this improvement increases in magnitude, reflecting the heightened influence of the filtering model. We see that this improvement increases with the error term $\lambda_{s}$, which scales the quality scores of tactics by their predicted probability of success, as was intended.
Table 5: Tactic success rate per node for miniF2F-valid (Mean $\pm$ Standard Error), where $\lambda_{s}$ controls the error weight of quality score in 3D-Prover. No filtering gives 27.7% $\pm$ 0.2%. We see that 3D-Prover leads to fewer errors on average, which can be controlled by increasing $\lambda_{s}$.
$K$ Top-$K$ Random 3D-Prover ($\lambda_{s}=0.1$) 3D-Prover ($\lambda_{s}=0.5$)
@@
Diversity
To measure diversity, we examine the percentage of successful tactics which result in a new proof path. We restrict to successful tactics to account for the discrepancy in success rate between approaches. We observe 3D-Prover gives more unique subgoals per successful tactic, which is noteworthy given the higher rate of successful tactics from 3D-Prover overall (Table 5). As intended, increasing $\theta$ gives further improvements under these metrics. This demonstrates that our approach is effective at avoiding redundant tactics, instead selecting tactics which yield more unique proof paths. Appendix F provides additional analysis, further supporting our claim of improved diversity.
Table 6: Percentage of successful tactics per node resulting in unique subgoal(s) over miniF2F-valid (Mean $\pm$ Standard Error). No filtering gives 67.8% $\pm$ 0.3%. We observe 3D-Prover results in more unique subgoals per tactic, leading to a more diverse set of proof paths, with larger $\theta$ controlling this.
$K$ Top-$K$ Random 3D-Prover ($\theta=1$) 3D-Prover ($\theta=4$)
8 85.3 $\pm$ 0.1 89.9 $\pm$ 0.1 90.1 $\pm$ 0.1 91.1 $\pm$ 0.1
16 77.5 $\pm$ 0.1 84.1 $\pm$ 0.1 84.9 $\pm$ 0.1 85.5 $\pm$ 0.1
32 72.3 $\pm$ 0.2 76.3 $\pm$ 0.2 76.9 $\pm$ 0.2 77.5 $\pm$ 0.2
Execution Time
Table 7 shows the execution time for tactics over miniF2F-valid transitions. Again we see that 3D-Prover outperforms the baselines, with the improvement increasing with more filtering. Increasing the time weight $\lambda_{\tau}$ results in further reductions to the average execution time, demonstrating the accuracy of the predictions, and that they can directly result in faster tactics when filtering. It has been noted that preferring faster tactics can prevent the excessive application of powerful automation tactics such as simp (Lample et al., 2022, Appendix E). As these generally take longer to run, using faster tactics can help models learn underlying proof arguments which are often hidden by the automation. It can also greatly reduce the number of timeout errors.
[…]
$K$ Top-$K$ Random 3D-Prover ($\lambda_{\tau}=0.1$) 3D-Prover ($\lambda_{\tau}=1.0$)
8 206 $\pm$ 0.8 198 $\pm$ 0.9 155 $\pm$ 0.5 136 $\pm$ 0.5
16 220 $\pm$ 0.8 218 $\pm$ 0.9 176 $\pm$ 0.6 152 $\pm$ 0.5
32 224 $\pm$ 0.8 215 $\pm$ 0.8 191 $\pm$ 0.7 181 $\pm$ 0.6
4 Conclusion
Limitations
Our main limitation was the scale of experiments we were able to run, with other results often requiring thousands of hours of train time using hundreds of provers and larger models (Lample et al., 2022; Polu et al., 2022; Xin et al., 2024). Given the large computational cost of evaluations (as we outline in Appendix I), we were only able to test InternLM2.5 up to Pass@2 for BFS and Pass@6 for Critic Guided, and ReProver up to Pass@4 on miniF2F-test, with a hyperparameter analysis of 15 runs on miniF2F-valid (Table C). We could only evaluate over the large LeanDojo benchmark (Appendix D) for single run with each approach.
Summary
We demonstrate the feasibility of learning proof system dynamics, where we generate tactic representations reflecting the response of the proof environment. We then leverage these with 3D-Prover, which filters candidate tactics to diverse and high quality subsets based on their likely outcome. We evaluate 3D-Prover by augmenting popular proving LLMs on the standard miniF2F and LeanDojo benchmarks, where we find an improvement in the overall proof success rate, particularly for deeper searches. Our ablation studies confirm the utility of our tactic representations, enabling the selection of tactics with improved success rates, diversity, and/or execution time. By effectively pruning the search space, 3D-Prover is a step towards enabling deeper automated proofs.
5 Acknowledgements
We acknowledge Defence Science and Technology Group (DSTG) for their support in this project.
Appendix A Proof Search Setup
We first define the space of goals $\mathcal{S}$, tactics $\mathcal{T}$ and failures $\mathcal{F}$. For our purposes, these all contain arbitrary strings, with the goal being a formal proposition, the tactic a command and the failure an error message. We then define the output space as ${\mathcal{O}}:=\mathcal{P}(\mathcal{S})\cup\mathcal{F}$. A proof tree is a DAG $G=(V,E)$ where $V\subset\mathcal{S}$ is the set of goals and $E$ the edges between them. A proof attempt for a goal $g_{0}$ first initialises the proof tree with $V=\{g_{0}\},E=\emptyset$. The search policy $\pi_{S}:G\times V\rightarrow\mathbb{R}^{+}$ is a distribution over goals given a proof tree, being used to select a goal $g$ to expand. The tactic policy $\pi_{T}:\mathcal{S}\times\mathcal{T}\to\mathbb{R}^{+}$ is a distribution over tactics given a goal, where $N\in\mathbb{N}$ tactics are sampled to give tactics $\{t_{i}\}_{i=1}^{N}\subset\mathcal{T}$. The goal, tactic pairs $(g,t_{i})$ are then passed to the environment $\mathcal{E}:\mathcal{S}\times\mathcal{T}\to\mathcal{O}$. For each pair, after $\tau_{i}\in\mathbb{R}$ seconds, it returns either a new set of goals $g_{i}^{\prime}\subset\mathcal{S}$ or an error, $e_{i}\in\mathcal{F}$. We define this response as the output $o_{i}\in\mathcal{O}$. We further define the status $s_{i}\in\{0,1\}$ as 0 if $o_{i}\in\mathcal{F}$, $1$ if $o_{i}\in\mathcal{P}(\mathcal{S})$ and the transition as the tuple $(g,t_{i},s_{i},\tau_{i},o_{i})$. The proof tree is then updated with $G=G\cup g_{i}^{\prime}$ for all $g_{i}^{\prime}$, and the associated transitions are added as edges to $E$. This is repeated until a proof is found, or a budget is exhausted. A proof of $g$ is found when $\mathcal{E}(g,t_{i})=\emptyset$ for any $t_{i}$, or if all $\{g_{i}^{\prime}\}$ are proven for $\mathcal{E}(g,t_{i})=\{g_{i}^{\prime}\}\subset\mathcal{S}$. The result of a proof attempt is then the set of transitions $\{(g_{k},t_{ki},s_{ki},\tau_{ki},o_{ki})\}$ for all selected goals $g_{k}$ and their expanded tactics $t_{i}$. For simplicity, we drop the indices to denote the set of transitions as $\{(g,t,s,\tau,o)\}$.
Appendix B Pass@k
Table 8: Percentage of proofs found after four attempts (Pass@4) with ReProver on miniF2F-test, with $K$ tactics selected per node.
$K$ Random 3D-Prover Gain
8 25.7 28.6 +11.3%
16 30.2 31.0 +2.6%
32 29.8 29.8 +0.0%
Table 9: Pass@$k$ rates for proof attempts on miniF2F-test
3D-Prover Random
Pass@$k$ $K=8$ $K=16$ $K=32$ $K=8$ $K=16$ $K=32$
1 24.9 27.8 28.6 18.0 21.2 28.1
2 26.1 29.4 29.0 22.9 28.6 29.0
3 26.5 29.8 29.8 24.9 29.4 29.8
4 28.6 31.0 29.8 25.7 30.2 29.8
Table 8 summarises the Pass@4 results for ReProver over miniF2F-test, which is the number of proofs found at least once over four attempts, with Table 9 showing the Pass@$k$ up to $k=4$. We compare 3D-Prover to the Random baseline, taking the same four runs from Table 2, where $\lambda_{s}=\lambda_{\tau}=0$, $\theta=1$. With Top-$K$ being deterministic, the Pass@$k$ rate is the same as the Pass@1 rate. Given several attempts, $K=16$ appears to provide a good tradeoff between breadth and depth, performing the best overall. 3D-Prover maintains a large improvement for $K=8$, with a modest improvement for $K=16$.
Appendix C Proof success rate over hyperparameters
Table 10: Pass@1 results on miniF2F-valid, over different hyperparameter configurations for 3D-Prover with ReProver.
($\lambda_{s}$, $\lambda_{\tau}$, $\theta$)
8 25.0 25.0 25.8 22.5 23.8
16 29.1 28.7 27.9 27.0 26.6
32 28.7 28.3 28.7 27.9 27.0
Appendix D Evaluation on LeanDojo Benchmark
We ran an additional experiment on the LeanDojo Novel Premises Yang et al. (2023) benchmark testing 3D-Prover on a larger dataset. This dataset has 2000 evaluation proofs in comparison to the 244 from miniF2F-valid and miniF2F-test, allowing us to evaluate the performance of 3D-Prover on a larger scale.
We trained a transition model from a single ReProver attempt on LeanDojo Novel Premises, before evaluating 3D-Prover using ReProver with the methodology in Section 4(a). We set KK=32 for 3D-Prover, and compare to the model with No Filtering (i.e. KK=64), and Top-KK=32. We further examine the distribution of proof lengths found from this experiment. To account for different proofs of the same goal, we adjust proof lengths to be the shortest found from any attempt ( e.g. if 3D-Prover finds a proof of length 10, which was found in 3 steps by No Filtering, we count it as length 3). Hence, all proof lengths reported are the shortest found by any method. We report the number of proofs found by each approach, organised by the proof length in Table 11.
Table 11: Number of Proofs found on LeanDojo Novel Premises, sorted by proof length.
Proof Length 3D-Prover ($K$ = 32) Top-$K$ ($K$ = 32) No Filtering ($K$ = 64)
1 236 233 237
2 167 162 174
3 134 126 131
4 60 60 54
5 40 39 24
6 7 6 2
7 2 0 0
Total 646 626 622
Pass@1 32.3% 31.3% 31.1%
3D-Prover obtains a relative improvement of 3.2% over Top-$K$, and a 3.9% relative improvement over No Filtering in terms of the number of proofs found, comparable to the same performance gain for $K=32$ found in miniF2F-valid (Table  2). We see that 3D-Prover finds deeper proofs, while maintaining a high proof success rate for shallower proofs, unlike Top-$K$. The No Filtering approach, as expected, finds the most shallow proofs, however quickly drops off in performance for deeper proofs. We also note that 3D-Prover found the 2 longest proofs of length 7, with neither baseline finding any.
Appendix E Embedding Discussion
Embedding Comparison
Embedding Objective
As outlined in Section 2, we train our embeddings to be reflective of the tactic semantics across all three components of Status, Time and Output. Hence 3D-Prover, which selects diverse embeddings, may lead to tactics predicted to have errors, where the errors are diverse in terms of their predicted message. The hyperparameter $\lambda_{s}$ can alleviate this by weighting the scores based on their likelihood of success. From our experiments (Table 10), there is not necessarily a benefit to Pass@1 by filtering out strongly based on the predicted error likelihood. To speculate, the error prediction, although quite good, is imperfect with false negatives (Table 1). This can lead to potentially useful tactics being ignored if the error prediction is overly trusted, even though there is a higher tactic success rate overall as in Table 5. Given these prediction errors, it may be the case that selecting goals which are predicted to lead to (diverse) errors may be preferable, given the possibility they result in successful new subgoals. These subgoals may be be quite different from those previously selected, as they are mispredicted, so are clearly outside the space of tactics where the transition model is confident about the outcome. Further analysis could be worthwhile to investigate this. An embedding architecture trained only on successful tactics could be used, however given the high error rate of tactics, this would ignore a large proportion of the transition data.
Appendix F Additional Diversity Analysis
To further examine diversity, we first look at the percentage of unique environment responses to tactics executed per node, including responses with unique errors (Table 12), using the same ReProver setup in 4(a). As it is difficult to select tactics guaranteed to be successful (see Table 5), we would expect a good exploratory policy to select tactics which result in more varied outputs (both errors and successes alike), so as to better explore the space. We also examine the degree of cosine similarity across unique subgoals (Table 13), using a simple text embedding model (all-MiniLM-L6-v2). This more precisely quantifies the diversity in the contents of the subgoals. We do this to account for simple changes (such as variable renaming), which would not be differentiated under a simple check for uniqueness, and so allows us to examine how varied the contents of the resulting subgoals are.
In both cases, we see that 3D-Prover results in more diverse responses. As intended, increasing $\theta$ results in further improvements to diversity under both metrics. The increased diversity in subgoal content (Table 13) is strengthened by the fact that 3D-Prover also gives more unique subgoals on average (Table 6), suggesting that our approach yields more unique subgoals, and that these subgoals are more varied in their contents.
Table 12: Percentage of unique environment responses per node in miniF2F-valid (Mean $\pm$ Standard Error). Unique defines either syntactically distinct error messages or responses including at least one previously unseen subgoal. No filtering results in 63.3% $\pm$ 0.2%. We see that 3D-Prover gives a higher diversity of environment responses, increasing with the diversity parameter $\theta$.